OpenStudy (anonymous):

3. Solve the equation and write the solution set. |4x - 8|- 3x = 6x + 2

OpenStudy (anonymous):

|4x-8| = 6x + 2 + 3x |4x-8| = 9x + 2 Next, find the critical points so lets say: +(4x-8) = 9x + 2 -(4x-8) = 9x + 2 x = -2, 6/13